OpenStudy (anonymous):

3over y =y over 6y-15

OpenStudy (anonymous):

Cross mulitiply

OpenStudy (anonymous):

15/16=Y

OpenStudy (anonymous):

\[\frac{3}{y}=\frac{y}{6y-15}\] \[3(6y-15)=y^2\] \[18y-45=y^2\] \[y^2-18y+48=0\]

OpenStudy (anonymous):

i meant 45 not 48!

OpenStudy (anonymous):

\[y^2-18y+45=0\]

OpenStudy (anonymous):

\[(y-3)(y-15)=0\] \[y=3\] or \[y=15\]

OpenStudy (anonymous):

and check to make sure that neither of these solutions will make the denominator 0. i.e. you must check your answers